Use the data table to answer the questions please. Thank you in advance! :)
Gas Laws Lab: Part 1: Collecting the gas in a balloon Calculations 1) Calculate the actual number of CO2 gas moles (n) in the balloon using the ideal gas Law PV= nRT (Show all your work) Assume: P pressure of CO2 in atm R 0.0821 L atm/ mol K T measured temperature of the room in K V volume of inflated balloon n number of moles of CO 2) Calculate the actual yield (in grams) of the CO2 gas from the actual number of gas moles and the molar mass of CO2 (Show all your work) Actual yield CO2 in grams 3) Calculate the theoretical yield (in grams) of CO2 gas produced from your starting mass of sodium bicarbonate, the limiting reactant. (Show all your work) Theoretical yield CO2 in grams = 4) Calculate the Percent Yield CO2 for the reaction using the formula: % Yield = Actual Yield (Show all your work) x 100 Theoretical Yield % Yield CO2 =
Gas Laws Lab: Part 1: Collecting the gas in a balloon Data table: Mass of sodium bicarbonate used (you aimed for 2.4 grams) 2.49 150 mL 29 9cm Volume of acetic acid (you aimed for 150 mL) Circumference of the balloon (cm) Temperature (°C) 20.50 Convert the temperature into Kelvin K °C 273 293.SKLC Atmospheric pressure (Pr) (atm) L028 03 Partial pressure of the water (PH2O) (see appendix B in your textbook) Pressure of CO2 (pco2) (atm) Pcoa = PT -PH20 그647- Determine the radius (r) of the inflated balloon in cm. 4.76 cm r C 2 TT C = circumference ㅠ 3D3.14 Determine volume (V) of inflated balloon in cm3 (assume a spherical shape) 451.4451 V= 4TT r 3 3 Convert the volume in cm3 to Liters (L) 1 cm3 = 1 mL 451cm3 at thermal equilibrium room temperature = CO2 gas temperature
Answer #1

1.

P=1.03

V=451ml

T=293.5K

R=0.821L

n=?

We know that pv=nRt

n=PV/RT =1.03*451/0.821*293.5 =1.92

The data is insufficient for ques 2 & 4.....